Will you please help me with this problem?
Write a system of two equations in two unknowns. Solve the system by using the addition method. « Mike paid $2.98 for three doughnuts and two coffees. The next day, he paid $1.64 for two doughnuts and one coffee. How much does it cost for one doughnut and one coffee?
Thank you!

Answer by stanbon(48568)

Write a system of two equations in two unknowns. Solve the system by using the addition method. « Mike paid $2.98 for three doughnuts and two coffees. The next day, he paid $1.64 for two doughnuts and one coffee. How much does it cost for one doughnut and one coffee?
-------------------------
1st: 3d+2c=2.98
2nd: 2d+ c=1.64
----------------
Multiply 2nd by 2 to get:
3rd: 4d+2c=3.28
-------------
Subtract 1st from 3rd to get:
d=0.30 ( Cost of one doughnut)
----------------
Substitute into 2nd to solve for c, as follows:
2(0.30)+c=1.64
c=1.04 (cost of one coffee)
Cheers,
Stan H.